# Kővári-Sós-Turán theorem for hereditary families

### v1

@@ 1 Introduction
Structural graph theory. Given a graph $H$, or a family of graphs $\mathcal{H}$, we are interested in the family $\mathcal{G}$ defined as the family of all graphs containing no induced copy of a member of $\mathcal{H}$. In this case, let us write $\mathrm{ex}^{*}(n,\mathcal{H},s):=\mathrm{ex}_{\mathcal{G}}(n,s)$, and simply $\mathrm{ex}^{*}(n,H,s)$ if =$\mathcal{H}=\{H\}$. It was proved by Kühn and Osthus [37] that if $\mathcal{H}$ is the family of all subdivisions of a fixed graph $H$, then $\mathrm{ex}^{*}(n,\mathcal{H},s)=O_{s}(n)$. The constant hidden in the $O_{s}(.)$ notation was recently improved by Du, Girão, Hunter, McCarthy, and Scott [12]. Bonamy et al. [6] showed that $\mathrm{ex}^{*}(n,P_{t},s)=s^{O_{t}(1)}n$ and $\mathrm{ex}^{*}(n,\mathcal{C}_{\geq t},s)=s^{O_{t}(1)}n$, where $P_{t}$ is the path of length $t$ and $\mathcal{C}_{\geq t}$ is the family of cycles of length at least $t$. A common strengthening of the previous results is proved independently by Girão and Hunter [25] and Bourneuf, Bucić, Cook and Davies [8]: if $\mathcal{H}$ is the family of subdivisions of a graph $H$, then $\mathrm{ex}^{*}(n,\mathcal{H},s)\leq s^{O_{H}(1)}n$. As another generalization of the result on paths, Scott, Seymour and Spirkl [48], improving a previous unpublished result of Rödl, proved that for every tree $T$, one has $\mathrm{ex}^{*}(n,T,s)=s^{O_{T}(1)}n$, where the exponent of $s$ has the order of magnitude $|T|^{\Omega(|T|)}$.
@@
Chan and Har-Peled [10] considered incidence graphs of $n$ points and $n$ pseudo-discs in the plane. A family of simple closed Jordan regions is a family of pseudo-discs if the boundary of any two intersect in at most two points. They proved that if such an incidence graph is $K_{s,s}$-free, then it has $O_{s}(n\log\log n)$ edges. This was strengthened by Keller and Smorodinsky [31] who proved that if $\mathcal{A}$ and $\mathcal{B}$ are families of $n$ pseudo-disks, then the bipartite intersection graph of $\mathcal{A}$ and $\mathcal{B}$ has at most $O(s^{6}n)$ edges, assuming it is $K_{s,s}$-free. We highlight that by a result Keszegh [32], such graphs contain no induced proper subdvision of a non-planar graph, in particular no induced proper subdivision of $K_{5}$. Hence, one immediately gets the bound $s^{O(1)}n$ by [8, 25]. The study of the family $\mathcal{G}_{\text{box}}$ of incidence graphs of points and axis-parallel boxes in $\mathbb{R}^{d}$ was initiated by Basit, Chernikov, Starchenko, Tao, and Tran [5]. Chan and Har-Peled [10] determined the optimal bound $\mathrm{ex}_{\mathcal{G}_{\text{box}}}(n,s)=O_{s}(n(\frac{\log n}{\log\log n})^{d-1})$. Furthermore, Tomon and Zakharov [55] study the closely related family of intersection graphs of axis-parallel boxes in $\mathbb{R}^{d}$, see also [31].
@@
As a high-dimensional generalization of the Szemerédi-Trotter theorem [53], Chazelle [11] proposed to study $K_{s,s}$-free incidence graphs of points and hyperplanes in $\mathbb{R}^{d}$. In this context, forbidding $K_{s,s}$ in the incidence graph is a natural nondegeneracy condition, since otherwise all points may lie on a line which is contained in all hyperplanes, in which no nontrivial upper bound on the number of incidences can be given. For $d\geq 3$, the currently best known upper bound on the number of edges in a $K_{s,s}$-free incidence graph of $n$ points and $n$ hyperplanes is $O_{d,s}(n^{2-2/(d+1)})$, proved by Apfelbaum and Sharir [4], while the best known lower bound is due to Sudakov and Tomon [51]. In [43], Milojević, Sudakov, and Tomon show that over any field $\mathbb{F}$, a $K_{s,s}$-free incidence graph of $n$ points and $n$ hyperplanes has at most $O_{d,s}(n^{2-\frac{1}{\lceil(d+1)/2\rceil}})$ edges, and this bound is the best possible. The proof of this result proceeds by considering $\mathrm{ex}^{*}(n,\mathcal{H},s)$ for a simple finite family of bipartite graphs $\mathcal{H}$ avoided by point-hyperplane incidence graphs. In another direction, Fox, Pach, Scheffer, Suk and Zahl [21] proved that the bound of Apfelbaum and Sharir [4] also holds (up to an $o(1)$ error term in the exponent) for $K_{s,s}$-free semialgebraic graphs in dimension $d$ as well (we refer the interested reader to [21] for precise definitions). One of their key tools is a bound on the number of edges in a $K_{s,s}$-free graph of VC-dimension at most $d$. Here, a graph has VC-dimension at most $d$ if it contains no set $A$ of $d+1$ vertices such that for every $X\subset A$ there is a vertex joined to every vertex in $X$, but no vertex in $A\setminus X$. In [21], it is proved that a $K_{s,s}$-free graph of VC-dimension at most $d$ has at most $O_{d,s}(n^{2-1/d})$ edges, see also [31] for an alternative proof. For $d\geq 3$, this was improved to $o(n^{2-1/d})$ by Janzer and Pohoata [29].
@@ 1.1 Main results
In this section, we present our results. Recall that $\mathrm{ex}^{*}(n,H,s)$ denotes the maximum number of edges in an $n$-vertex graph containing no $K_{s,s}$ and no induced copy of $H$. We consider $\mathrm{ex}^{*}(n,H,s)$ for various bipartite graphs $H$. Note that the case when $H$ is not bipartite is not very interesting: as every graph contains a bipartite subgraph with at least half of the edges, we have $\mathrm{ex}^{*}(n,H,s)=\Theta(\mathrm{ex}(n,K_{s,s}))$. In case $H$ is bipartite, we believe that for sufficiently large $s$, $\mathrm{ex}^{*}(n,H,s)$ is close to the extremal number of $H$. In particular, we propose the following conjecture.
@@
For every bipartite graph $H$,
@@
Next, we consider trees. It is an easy exercise to show that if $T$ is a tree, then $\mathrm{ex}(n,T)\leq|T|n$. As proved by Scott, Seymour, and Spirkl [48], $\mathrm{ex}^{*}(n,T,s)$ also grows linearly as a function of $n$, and furthermore as a polynomial function of $s$. The exponent of $s$ in their result is super-exponential in $|T|$, it is of the order $|T|^{\Omega(|T|)}$. One of our main results improves this to linear, i.e., $\mathrm{ex}^{*}(n,T,s)=s^{O(|T|)}n$. This is optimal up to the constant term hidden by $O(.)$ for every tree $T$.
@@ 1.2 Applications
Our results and techniques have several additional application which we discuss in this subsection.
Kühn and Osthus [36] proved that for every $k$, every graph of sufficiently large average degree contains a $C_{4}$-free subgraph of average degree at least $k$ (see also [44] for improved bounds). McCarty [42] established a natural induced variant of this result: for every $k$ and sufficiently large $s$, there exists a smallest number $g(k,s)$ such that if a $K_{s,s}$-free graph $G$ has average degree at least $g(k,s)$, then $G$ contains a $C_{4}$-free induced subgraph of average degree at least $k$. Quantitative bounds are first proved by Du, Girão, Hunter, McCarty, and Scott [12], who show that $g(k,s)\leq k^{O(s^{3})}$, while Girão and Hunter [25] prove that $g(k,s)\leq s^{O(k^{4})}$. In [25], the lower bound $g(k,s)\geq s^{\Omega(k^{2})}$ is also established for $s$ sufficiently large with respect to $k$. Our results can be used to show that this lower bound is tight and $g(k,s)\leq s^{O(k^{2})}$.
@@
A well known lopsided weakening of the Erdős-Hajnal conjecture, proved by Fox and Sudakov [22], says that every induced $H$-free graph on $n$ vertices contains either a complete bipartite graph with vertex classes of size $n^{c}$, or an independent set of size at least $n^{c}$, for some $c=c(H)>0$. Their proof implies that one can take $c(H)=\Omega(1/|V(H)|^{3})$. We improve this in case $H$ is bipartite to $c(H)=\Omega(1/|V(H)|)$, which is optimal as can be shown by considering appropriate random graphs. We also remark that a recent result of Nguyen, Scott, and Seymour [45] proves the Erdős-Hajnal conjecture in the case one forbids bi-induced copies of a bipartite graph $H$, i.e. we forbid every induced copy of those graphs we get by possibly adding edges to the two vertex classes of $H$.
@@ 2.1 Embedding the vertices of degree at most $k$
We embed the vertices $b_{1},b_{2},\dots,b_{|B|}$ inductively. Assume that we already defined the images $\phi(b_{1}),\dots,\phi(b_{i})$. A vertex $v\in V(G)\backslash(\{\phi(b_{1}),\dots,\phi(b_{i})\}\cup I)$ is a candidate for $b_{j}$ if it is not adjacent to $\phi(b_{1}),\dots,\phi(b_{i})$, and satisfies $N(v)\cap I=\phi(N_{H}(b_{j}))$. The first condition serves to ensure that the embedding of $B$ forms an independent set, while the second one guarantees that the edges between $\phi(A)$ and $\phi(B)$ correspond precisely to the edges of $H$. We show by induction on $i=0,1,\dots,|B|$ that one can embed the vertices $b_{1},\dots,b_{i}$ such that for each $j>i$, there are at least $(4|B|s)^{|B|-i}$ candidates for $b_{j}$. When $i=0$, there are at least $(4|B|s)^{|B|}$ candidates for each $b_{j}$, since $I$ is a rich independent set and $b_{j}$ has at most $k$ neighbours. Having defined $\phi(b_{1}),\dots,\phi(b_{i})$, our goal is to define $\phi(b_{i+1})$ such that the sets of candidates not decreased too much. More precisely, if we denote the set of candidates for $b_{j}$ by $U_{j}$, our goal is to find a vertex $v\in U_{i+1}$ for which $|U_{j}\backslash N(v)|\geq|U_{j}|/2s$ for all $j\geq i+2$. If we find such a vertex $v$, then we take $\phi(b_{i+1}):=v$.
@@ 2.2 Finding a rich independent set
Let us now give an imprecise explanation of how these notions will come into play in our proof. We consider the hypergraph $\mathcal{H}$ which consists of many independent sets of a given size inside the set $X\subseteq V(G)$. The first step is to turn this hypergraph into a $(\varepsilon,\delta)$-superspread hypergraph by cleaning it. Then, we define a collection of “bad” tuples $\mathfrak{B}$ by saying that $(v_{1},\dots,v_{\ell})$ is bad if $N(v_{\ell})$ contains too many vertices from the set $S=N(v_{1},\dots,v_{i})\backslash\bigcup_{j=i+1}^{\ell-1}N(v_{j})$, for some fixed index $i$. Thus, for any fixed $v_{1},\dots,v_{\ell-1}$ for which the set $S$ is not too small, Claim 2.2 shows that there are at most $s$ vertices $v_{\ell}$ making the $\ell$-tuple $(v_{1},\dots,v_{\ell})$ bad. Then, we show that there exists an independent set avoiding all bad tuples, which turns out to be sufficient to show for the independent set to be rich.
@@
    Lemma 2.6 shows that if $\mathcal{H}_{0}$ is not $(\varepsilon,\delta)$-superspread, one can eliminate one of the parts $V_{i}$ and obtain a $(r-1)$-partite $(r-1)$-uniform hypergraph $\mathcal{H}_{1}$ with at least $\frac{\varepsilon\delta}{r2^{r}}e(\mathcal{H}_{0})$ edges. In general, this procedure can be repeated, thus obtaining a sequence of hypergraphs $\mathcal{H}_{0},\mathcal{H}_{1},\dots,\mathcal{H}_{m}$, where $\mathcal{H}_{i}$ is a $(r-i)$-partite $(r-i)$-uniform graph with
@@
    The last equality is the consequence of double counting, as the sum $\sum\deg_{\mathcal{H}}(v_{1},\dots,v_{\ell-1})$ counts the number of pairs $(e,(v_{1},\dots,v_{\ell-1}))$ for which the vertices $v_{1},\dots,v_{\ell-1}$ belong to $e$ and $e\in E(\mathcal{H})$. Since every edge $e$ of cardinality $t$ contain exactly $(\ell-1)!\binom{t}{\ell-1}$ ordered $(\ell-1)$-tuples, the equality follows. Since $\delta s(\ell-1)!\binom{t}{\ell-1}e(\mathcal{H})<\delta st!e(\mathcal{H})$, we conclude that the number of bad edges is at most $\delta st!e(\mathcal{H})$, finishing the proof. ∎
@@
    Set $r:=2a$. The main idea is to apply Proposition 2.5 to the $r$-uniform hypergraph, whose edges are the independent sets of size $2a$ in $X$. In order to do this, we need to verify that there are at least $C_{r}(\varepsilon\delta)^{-r}|X|^{a-1}$ independent sets of size $r$ in $X$, where $C_{r}=(r)^{r}2^{r^{2}}$.

### v2

@@ 1 Introduction
Structural graph theory. Given a graph $H$, or a family of graphs $\mathcal{H}$, we are interested in the family $\mathcal{G}$ defined as the family of all graphs containing no induced copy of a member of $\mathcal{H}$. In this case, let us write $\mathrm{ex}^{*}(n,\mathcal{H},s):=\mathrm{ex}_{\mathcal{G}}(n,s)$, and simply $\mathrm{ex}^{*}(n,H,s)$ if =$\mathcal{H}=\{H\}$. It was proved by Kühn and Osthus [38] that if $\mathcal{H}$ is the family of all subdivisions of a fixed graph $H$, then $\mathrm{ex}^{*}(n,\mathcal{H},s)=O_{s}(n)$. The constant hidden in the $O_{s}(.)$ notation was recently improved by Du, Girão, Hunter, McCarty and Scott [13]. Bonamy et al. [7] showed that $\mathrm{ex}^{*}(n,P_{t},s)=s^{O_{t}(1)}n$ and $\mathrm{ex}^{*}(n,\mathcal{C}_{\geq t},s)=s^{O_{t}(1)}n$, where $P_{t}$ is the path of length $t$ and $\mathcal{C}_{\geq t}$ is the family of cycles of length at least $t$. A common strengthening of the previous results is proved independently by Girão and Hunter [26] and Bourneuf, Bucić, Cook and Davies [9]: if $\mathcal{H}$ is the family of subdivisions of a graph $H$, then $\mathrm{ex}^{*}(n,\mathcal{H},s)\leq s^{O_{H}(1)}n$. As another generalization of the result on paths, Scott, Seymour and Spirkl [49], improving a previous unpublished result of Rödl, proved that for every tree $T$, one has $\mathrm{ex}^{*}(n,T,s)=s^{O_{T}(1)}n$, where the exponent of $s$ has the order of magnitude $|T|^{\Omega(|T|)}$.
@@
Chan and Har-Peled [11] considered incidence graphs of $n$ points and $n$ pseudo-discs in the plane. A family of simple closed Jordan regions is a family of pseudo-discs if the boundary of any two intersect in at most two points. They proved that if such an incidence graph is $K_{s,s}$-free, then it has $O_{s}(n\log\log n)$ edges. This was strengthened by Keller and Smorodinsky [32] who proved that if $\mathcal{A}$ and $\mathcal{B}$ are families of $n$ pseudo-disks, then the bipartite intersection graph of $\mathcal{A}$ and $\mathcal{B}$ has at most $O(s^{6}n)$ edges, assuming it is $K_{s,s}$-free. We highlight that by a result Keszegh [33], such graphs contain no induced proper subdivision of a non-planar graph, in particular no induced proper subdivision of $K_{5}$. Hence, one immediately gets the bound $s^{O(1)}n$ by [9, 26]. The study of the family $\mathcal{G}_{\text{box}}$ of incidence graphs of points and axis-parallel boxes in $\mathbb{R}^{d}$ was initiated by Basit, Chernikov, Starchenko, Tao and Tran [6]. Chan and Har-Peled [11] determined the optimal bound $\mathrm{ex}_{\mathcal{G}_{\text{box}}}(n,s)=O_{s}(n(\frac{\log n}{\log\log n})^{d-1})$. Furthermore, Tomon and Zakharov [57] study the closely related family of intersection graphs of axis-parallel boxes in $\mathbb{R}^{d}$, see also [32].
@@
As a high-dimensional generalization of the Szemerédi-Trotter theorem [55], Chazelle [12] proposed to study $K_{s,s}$-free incidence graphs of points and hyperplanes in $\mathbb{R}^{d}$. In this context, forbidding $K_{s,s}$ in the incidence graph is a natural nondegeneracy condition, since otherwise all points may lie on a line which is contained in all hyperplanes, in which case no nontrivial upper bound on the number of incidences can be given. For $d\geq 3$, the currently best known upper bound on the number of edges in a $K_{s,s}$-free incidence graph of $n$ points and $n$ hyperplanes is $O_{d,s}(n^{2-2/(d+1)})$, proved by Apfelbaum and Sharir [3], while the best known lower bound is due to Sudakov and Tomon [53]. In [44], Milojević, Sudakov and Tomon show that over any field $\mathbb{F}$, a $K_{s,s}$-free incidence graph of $n$ points and $n$ hyperplanes has at most $O_{d,s}(n^{2-\frac{1}{\lceil(d+1)/2\rceil}})$ edges, and this bound is the best possible. The proof of this result proceeds by considering $\mathrm{ex}^{*}(n,\mathcal{H},s)$ for a simple finite family of bipartite graphs $\mathcal{H}$ avoided by point-hyperplane incidence graphs. In another direction, Fox, Pach, Scheffer, Suk and Zahl [22] proved that the bound of Apfelbaum and Sharir [3] also holds (up to an $o(1)$ error term in the exponent) for $K_{s,s}$-free semialgebraic graphs in dimension $d$ as well (we refer the interested reader to [22] for precise definitions). One of their key tools is a bound on the number of edges in a $K_{s,s}$-free graph of VC-dimension at most $d$. Here, a graph has VC-dimension at most $d$ if it contains no set $A$ of $d+1$ vertices such that for every $X\subset A$ there is a vertex joined to every vertex in $X$, but no vertex in $A\setminus X$. In [22], it is proved that a $K_{s,s}$-free graph of VC-dimension at most $d$ has at most $O_{d,s}(n^{2-1/d})$ edges, see also [32] for an alternative proof. For $d\geq 3$, this was improved to $o(n^{2-1/d})$ by Janzer and Pohoata [31]. For a generalization of this result to an even wider class of graphs by Axenovich and Zimmermann, see [5].
@@ 1.1 Main results
In this section, we present our results. Recall that $\mathrm{ex}^{*}(n,H,s)$ denotes the maximum number of edges in an $n$-vertex graph containing no $K_{s,s}$ and no induced copy of $H$. We consider $\mathrm{ex}^{*}(n,H,s)$ for various bipartite graphs $H$. Note that the case when $H$ is not bipartite is not very interesting: as every graph contains a bipartite subgraph with at least half of the edges, we have $\mathrm{ex}^{*}(n,H,s)=\Theta(\mathrm{ex}(n,K_{s,s}))$. In the case where $H$ is bipartite, we believe that for sufficiently large $s$, $\mathrm{ex}^{*}(n,H,s)$ is close to the extremal number of $H$. In particular, we propose the following conjecture.
@@
For every connected bipartite graph $H$,
@@
Next, we consider trees. It is an easy exercise to show that if $T$ is a tree, then $\mathrm{ex}(n,T)\leq|T|n$. As proved by Scott, Seymour and Spirkl [49], $\mathrm{ex}^{*}(n,T,s)$ also grows linearly as a function of $n$, and furthermore as a polynomial function of $s$. The exponent of $s$ in their result is super-exponential in $|T|$, it is of the order $|T|^{\Omega(|T|)}$. One of our main results improves this to linear, i.e., $\mathrm{ex}^{*}(n,T,s)=s^{O(|T|)}n$. This is optimal up to the constant term hidden by $O(.)$ for every tree $T$.
@@ 1.2 Applications
Our results and techniques have several additional applications which we discuss in this subsection.
Kühn and Osthus [37] proved that for every $k$, every graph of sufficiently large average degree contains a $C_{4}$-free subgraph of average degree at least $k$ (see also [45] for improved bounds). McCarty [43] established a natural induced variant of this result: for every $k$ and sufficiently large $s$, there exists a smallest number $g(k,s)$ such that if a $K_{s,s}$-free graph $G$ has average degree at least $g(k,s)$, then $G$ contains a $C_{4}$-free induced subgraph of average degree at least $k$. Quantitative bounds are first proved by Du, Girão, Hunter, McCarty and Scott [13], who show that $g(k,s)\leq k^{O(s^{3})}$, while Girão and Hunter [26] prove that $g(k,s)\leq s^{O(k^{4})}$. In [26], the lower bound $g(k,s)\geq s^{\Omega(k^{2})}$ is also established for $s$ sufficiently large with respect to $k$. Our results can be used to show that this lower bound is tight and $g(k,s)\leq s^{O(k^{2})}$.
@@
A well known lopsided weakening of the Erdős-Hajnal conjecture, proved by Fox and Sudakov [23], says that every induced $H$-free graph on $n$ vertices contains either a complete bipartite graph with vertex classes of size $n^{c}$, or an independent set of size at least $n^{c}$, for some $c=c(H)>0$. Their proof implies that one can take $c(H)=\Omega(1/|V(H)|^{3})$. We improve this in case $H$ is bipartite to $c(H)=\Omega(1/|V(H)|)$, which is optimal as can be shown by considering appropriate random graphs. We also remark that a recent result of Nguyen, Scott and Seymour [46] proves the Erdős-Hajnal conjecture in the case one forbids bi-induced copies of a bipartite graph $H$, i.e. we forbid every induced copy of those graphs we get by possibly adding edges to the two vertex classes of $H$.
@@ 2.1 Embedding the vertices of degree at most $k$
We embed the vertices $b_{1},b_{2},\dots,b_{|B|}$ inductively. Assume that we already defined the images $\phi(b_{1}),\dots,\phi(b_{i})$. A vertex $v\in V(G)\backslash(\{\phi(b_{1}),\dots,\phi(b_{i})\}\cup I)$ is a candidate for $b_{j}$ if it is not adjacent to $\phi(b_{1}),\dots,\phi(b_{i})$, and satisfies $N(v)\cap I=\phi(N_{H}(b_{j}))$. The first condition serves to ensure that the embedding of $B$ forms an independent set, while the second one guarantees that the edges between $\phi(A)$ and $\phi(B)$ correspond precisely to the edges of $H$. We show by induction on $i=0,1,\dots,|B|$ that one can embed the vertices $b_{1},\dots,b_{i}$ such that for each $j>i$, there are at least $(4|B|s)^{|B|-i}$ candidates for $b_{j}$. When $i=0$, there are at least $(4|B|s)^{|B|}$ candidates for each $b_{j}$, since $I$ is a rich independent set and $b_{j}$ has at most $k$ neighbours. Having defined $\phi(b_{1}),\dots,\phi(b_{i})$, our goal is to define $\phi(b_{i+1})$ such that the sets of candidates have not decreased too much. More precisely, if we denote the set of candidates for $b_{j}$ by $U_{j}$, our goal is to find a vertex $v\in U_{i+1}$ for which $|U_{j}\backslash N(v)|\geq|U_{j}|/2s$ for all $j\geq i+2$. If we find such a vertex $v$, then we take $\phi(b_{i+1}):=v$.
@@ 2.2 Finding a rich independent set
Let us now give an imprecise explanation of how these notions will come into play in our proof. We consider the hypergraph $\mathcal{H}$ which consists of many independent sets of a given size inside the set $X\subseteq V(G)$. The first step is to turn this hypergraph into a $(\varepsilon,\delta)$-superspread hypergraph by cleaning it. Then, we define a collection of “bad” tuples $\mathfrak{B}$ by saying that $(v_{1},\dots,v_{\ell})$ is bad if $N(v_{\ell})$ contains too many vertices from the set $S=N(v_{1},\dots,v_{i})\backslash\bigcup_{j=i+1}^{\ell-1}N(v_{j})$, for some fixed index $i$. Thus, for any fixed $v_{1},\dots,v_{\ell-1}$ for which the set $S$ is not too small, Claim 2.2 shows that there are at most $s$ vertices $v_{\ell}$ making the $\ell$-tuple $(v_{1},\dots,v_{\ell})$ bad. Then, we show that there exists an independent set avoiding all bad tuples, which turns out to be sufficient for the independent set to be rich.
@@
    Lemma 2.6 shows that if $\mathcal{H}_{0}$ is not $(\varepsilon,\delta)$-superspread, one can eliminate one of the parts $V_{i}$ and obtain an $(r-1)$-partite $(r-1)$-uniform hypergraph $\mathcal{H}_{1}$ with at least $\frac{\varepsilon\delta}{r2^{r}}e(\mathcal{H}_{0})$ edges. In general, this procedure can be repeated, thus obtaining a sequence of hypergraphs $\mathcal{H}_{0},\mathcal{H}_{1},\dots,\mathcal{H}_{m}$, where $\mathcal{H}_{i}$ is an $(r-i)$-partite $(r-i)$-uniform hypergraph with
@@
    The last equality is the consequence of double counting, as the sum $\sum\deg_{\mathcal{H}}(v_{1},\dots,v_{\ell-1})$ counts the number of pairs $(e,(v_{1},\dots,v_{\ell-1}))$ for which the vertices $v_{1},\dots,v_{\ell-1}$ belong to $e$ and $e\in E(\mathcal{H})$. Since every edge $e$ of cardinality $t$ contains exactly $(\ell-1)!\binom{t}{\ell-1}$ ordered $(\ell-1)$-tuples, the equality follows. Since $\delta s(\ell-1)!\binom{t}{\ell-1}e(\mathcal{H})<\delta st!e(\mathcal{H})$, we conclude that the number of bad edges is at most $\delta st!e(\mathcal{H})$, finishing the proof. ∎
@@
    Set $r:=2a$. The main idea is to apply Proposition 2.5 to the $r$-uniform hypergraph, whose edges are the independent sets of size $2a$ in $X$. In order to do this, we need to verify that there are at least $C_{r}(\varepsilon\delta)^{-r}|X|^{a-1}$ independent sets of size $r$ in $X$, where $C_{r}=r^{r}2^{r^{2}}$.